Your father bought a car battery that has a terminal voltage of 12.6V for 75.0A, and 14.1V for 60.0A. What is the electromotive force and internal resistance of this car battery that your father bought?


1
Expert's answer
2022-03-02T10:49:14-0500

The Ohm's law says

"I_1=\frac{\cal E}{R_1+r}"

"I_2=\frac{\cal E}{R_2+r}"

or

"I_1=\frac{\cal E}{V_1/I_1+r}"

"I_2=\frac{\cal E}{V_2/I_2+r}"

We get

"75=\frac{\cal E}{12.6/75+r}"

"60=\frac{\cal E}{14.1/60+r}"


Roots:

"r=0.1\Omega, \;\cal E=20.1\:\rm V"